Common stock valuelong dashConstant growth   McCracken​ Roofing, Inc., common stock paid a dividend of ​$1.48 per share last year. The company expects earnings and dividends to grow at a rate of 6​% per year for the foreseeable future.  

a.  What required rate of return for this stock would result in a price per share of ​$28​?

b. If McCracken expects both earnings and dividends to grow at an annual rate of 12​%, what required rate of return would result in a price per share of ​$28​?
Answer #1
a. Required rate of return = (D0*(1+g)/P0)+g Where,
= (1.48*(1+0.06)/28.00)+0.06 D0 = $       1.48
= 11.60% g = 6%
P0 = $    28.00
b. Required rate of return = (D0*(1+g)/P0)+g Where,
= (1.48*(1+0.12)/28.00)+0.12 D0 = $       1.48
= 17.92% g = 12%
P0 = $    28.00
Answer #2

SOLUTION :


a.


Dividend paid last year, D0 = 1.48 ($)


Growth rate for dividend and earnings, g = 6% = 0.06 (constant)

=> 1 + g = 1.06


=> Divided for year 1, D1 = D0(1+g) = 1.48*1.06 = 1.5688 


Let required rate of return be r to get price of share at 28 ($)


As per CAPM :


Required rate of return, r : 

= D1/P + g 

= 1.5688/28 + 0.06

= 0.1160

= 11.60% (ANSWER).


b.


If g = 12% = 0.12

=> 1+g = 1.12

=> D1 = Do(1+g) = 1.48*1.12 = 1.6576 


So,


Required rate of return, r : 

= D1/P + g 

= 1.6576/28 + 0.12

= 0.1792

= 17.92% (ANSWER).

answered by: Tulsiram Garg